QUESTION 2

Babasiga.com, the online bookseller, wants to increase its total revenue. One strategy is to offer a 10% discount on every book it sells. Babasiga.com knows that its customers can be divided into two distinct groups according to their likely responses to the discount. The accompanying table shows how the two groups respond to the discount.

Group A

(sales per week)

Group B

(sales per week)


Volume of sales before the 10% discount

1.55 million

1.50 million

Volume of sales after the 10% discount

1.65 million

1.70 Million


a.Using the midpoint method, calculate the price elasticities of demand for group A and group B

(3Marks)





1
Expert's answer
2022-03-16T11:09:18-0400

"demand = \frac{(Q2 - Q1)}{ \frac{(Q2 + Q1)}{ 2}}= \frac{(1.65 - 1.55)}{ \frac{(1.65 + 1.55)}{ 2}}=0.0625"

"Price = \frac{(P2 - P1)}{ \frac{(P2 + P1)}{ 2}}= \frac{(1.70 - 1.50)}{ \frac{(1.70 + 1.50)}{ 2}}=0.125"


Dividend demand elasticity by price elasticity.


"Elasticity=\frac{0.0625}{0.125}=0.5"


Elasticity=50%